OpenStudy (tabdo1226):

Identify which table shows a direct variation. Two tables. For each table the column headings are x and y. Table 1 has column 1 values 3, 4.5, 7 and column 2 values 7.5, 11.25, 17.5. Table 2 has column 1 values 1.5, 8, 13 and column 2 values 3, 12, 19.5. A. both tables B. Table 1 only C. Table 2 only D. neither table

OpenStudy (tabdo1226):

@texaschic101

OpenStudy (tabdo1226):

do you know the answer @texaschic101

OpenStudy (texaschic101):

I might do this a little bit different then others...but direct variation reminds me of proportions. table 1 : 3/7.5 = 4.5/11.25 cross multiply (3)(11.25) = (7.5)(4.5) 33.75 = 33.75 (correct) This is direct variation table 2 : 1.5/3 = 8/12 cross multiply (1.5)(12) = (3)(8) 18 = 24 (incorrect) This is not direct variation

OpenStudy (tabdo1226):

thanks

OpenStudy (texaschic101):

sure thing :)

OpenStudy (tabdo1226):

so it's b table 1